Kara Less was applying her makeup when she drove into South's busy parking lot last Friday morning. Unaware that Lisa Ford was stopped in her lane 30 feet ahead, Kara rear-ended Lisa's rented Taurus. Kara's 1300-kg car was moving at 11 m/s and stopped in 0.14 seconds. 3pts each





a. Determine the momentum change of Kara's car.





b. Determine the impulse experienced by Kara's car.,





c. Determine the magnitude of the force experienced by Kara's car.

1
Expert's answer
2022-03-01T15:17:56-0500

a) The initial momentum is "p_i = 1300kg\cdot 11m/s = 14300kg\cdot m/s". The final momentum is "p_f = 0kg\cdot m/s". Thus, the change is:


"\Delta p = p_f - p_i\approx -14300kg\cdot m/s"

b) The impulse by definition is equal to the momentum change.

c) The force by the second Newton's law is given as follows:


"F = \dfrac{14000kg\cdot m/s}{0.14s} \approx 10^5N"